Grade 11General Physics

A body A is projected upwards with a vilocity of 9.8 m per second . The second body B is projected upwards with the same initial velocity but after 4 second . Both the bodies will meet after.
(A) 6 second
(B) 8 second
(C) 10 second
(D) 12 second

Aryan
8 Years agoGrade 11

1 Answer

Pranav singh
8 Years ago
Time  of flight for ball A =2s(v=u+at) ,time of flight for ball B-  0=9.8-9.8t => t=1  => time of flight =2s therefore time after which they will meet = 4+2+2=8 s